I'm not sure if this is the right category, but I'd really appreciate your help on this multiple choice question: " If 2+i and 2-i are the roots of the equation x^2 -4x +c= 0, what is the value of c?"
a) -5 b) 5 c) -4 d)4
Thanks!

Answer by stanbon(75887)

If 2+i and 2-i are the roots of the equation x^2 -4x +c= 0, what is the value of c?"
a) -5 b) 5 c) -4 d)4
=================================
If those are the roots then (x-(2-i))(x-(2+i)) must divide x^2-4x+c
-----------------
But that product is ((x-2)+i)((x-2)-i) = (x-2)^2+1 = x^2-4x+4+1
= x^2 -4x + 5
=========================
So c = 5
=============================
Cheers,
Stan H.